An electric dipole is made up of 2 particles having charges  1micro columb  mass 1 kg  and other with  charge _1 micro  columb  mass 1 kg separated by distance 1m it is in equilibrium  in a uniform  electric  field of 20×1000 volt per metre of dipole is deflected  through  an angle 2 degree  time taken by it to come again in equilibrium  is

Grade:12

2 Answers

Arun
5 years ago
Torque=Iα 
Torque=-pEsin? 
Iα=-pEsin? 
Contrasting this condition and the SHM condition 
mα=-kx 
rakish recurrence = √(k/m) 
k=pE 
=>k= 6x104 
Angular frequency  of the 2kg mass charge = √(6X?10?^4/2) = 100√3 
Angular frequency of the 1kg mass charge = √(6X?10?^4/1) = 100√6 
Sivabhanu
2 years ago
Torque =(I × α) ...1 ( alpha = acceleration ) And also Torque = PEsin θ .....2From 1 and 2 I ×α = PEsinθ. We know( I = mr^2 ) Here there are 2 bodies And r = L/2 Then I = 2×m×(L/2)^2 =(mL^2)/2 And α =( w^2) ×θ(mL^2)/2. ×(w^2)×θ = -q×E×L×θ. (Sinθ~θ.)W^2 =( -2qE/mL)W= √(2qE/mL)Time period = 2π/w = 2π×1/w = 2π√(mL/2qE)Dipole takes T/4 of time to come again equilibrium. t =(2π/4)√(mL/2qE)Subtituting values t =( π/2)√(1×1)/(2×10^-6 ×20×10^3) t = (π/2)×5 t.= 2.5π sec . Answer